# artificial intelligence CLI command-line interface

### v2

@@ 6.1 Limitations and Future Work
First, Innamark’s small embedding capacity is a weakness. The structure of an InnamarkTag, as shown in Fig. 1, is designed to enable optional compression that can help to increase the capacity. Work is in progress to increase the capacity further by using compression libraries or developing hybrid approaches that use non-printable characters.
Second, the proposed solution is based and tested on the Unicode standard and the UTF-8 scheme. Future research should consider the influence of other encoding schemes like UTF-16 and ISO-standardized Latin-1 (ISO 8859-1), and the potential impacts on the process of recoding to a small scheme like ASCII.
@@
Fifth, accessibility concerns may arise if a text with a secret message is read by screen readers or automatically translated into another language. Future research is needed to analyze how those accessibility tools handle specific Unicode characters.
@@ 7 Conclusion
We have designed and implemented Innamark, a blind and invisible information-hiding technique that can embed any byte-encoded sequence inside a cover text. Although several solutions for digital text watermarking and steganography have been published in recent years, existing approaches change the semantics or style of the cover text, increase the number of characters, or lack robustness against the output being copied into different applications. By encoding and mapping a secret message into our embedding alphabet of five Unicode whitespace characters, we can embed the information in the cover text by substituting all whitespace characters. The specified structure of our InnamarkTag has been designed to enable additional functionalities like compression, encryption, hashing, and error correction. The experimental evaluation shows strengths in imperceptibility and robustness, with limitations in embedding capacity based on a direct benchmark comparison with ten algorithms. Our method can help LLM operators fulfill regulations like the European Union’s AI Act [18] and can assist businesses in securing sensitive data before it is shared with external parties, especially if they are concerned about robustness to copying between applications. Work is in progress to increase the algorithm’s embedding capacity and enable the restoration of secret messages broken by text alterations.

### v3

@@ 6.1 Limitations and Future Work
First, Innamark’s small embedding capacity is a weakness. The structure of an InnamarkTag, as shown in Fig. 1, is designed to enable optional compression that can help to increase the capacity. Work is in progress to increase the capacity further by using compression libraries or developing hybrid approaches that use non-printable characters. Furthermore, the impact of using different InnamarkTag options needs to be analyzed to determine to what extent the error correction improves robustness or how hashing affects the embedding capacity.
Second, the proposed solution is based on and tested on the Unicode standard and the UTF-8 scheme. Future research should consider the influence of other encoding schemes like UTF-16 and ISO-standardized Latin-1 (ISO 8859-1), and the potential impacts on the process of recoding to a small scheme like ASCII.
@@
Fifth, accessibility concerns may arise if a text with a secret message is read by screen readers or automatically translated into another language. Future research is needed to analyze how those accessibility tools handle specific Unicode characters. This emphasizes the need for a deeper analysis of related text processing tools, such as machine learning-based natural language processing (NLP) units, minifiers, compressors, or deobfuscation tools.
@@ 7 Conclusion
We have designed and implemented Innamark, a blind and invisible information-hiding technique that can embed byte-encoded sequences inside a cover text. Although several solutions for digital text watermarking and steganography have been published in recent years, existing approaches change the semantics or style of the cover text, increase the number of characters, or lack robustness against the output being copied into different applications. By encoding and mapping a secret message into our embedding alphabet of five Unicode whitespace characters, we can embed the information in the cover text by substituting all whitespace characters. The specified structure of our InnamarkTag has been designed to enable additional functionalities like compression, encryption, hashing, and error correction. The experimental evaluation shows strengths in imperceptibility and robustness, with limitations in embedding capacity based on a direct benchmark comparison with ten algorithms. Our method can help LLM operators fulfill regulations like the European Union’s AI Act [18] and can assist businesses in securing sensitive data before it is shared with external parties, especially if they are concerned about robustness to copying between applications. Work is in progress to increase the algorithm’s embedding capacity and enable the restoration of secret messages broken by text alterations.